Material Engineering

Suppose 1 g of copper as Cu2+ is dissolved in 1000 g of water to produce an electrolyte. Calculate the electrode potential of the copper half-cell in this electrolyte. Assume T = 25oC. Express your answer in volts and to three significant figures. Tolerance ±0.2.
Answer #1

For getting one moles of Cu in 1000ml water gives rise to 1M of solution

Atomic mass of Cu= 63.54 gm

Concentration of 1 g= 1/63.54 M= 0.015738

From Nerst Equation

E= Eo-0.0592/n log (Cu+2]

Where n= number of electrons=2

E= 0.34-0.0592/2ln (1/0.0157)=0.34+ 0.0592/2 log(0.0157) =0.34-0.053=0.287V